Provide the missing data in the following table for a distributor of martial arts products: (Enter Turnover and ROI answe
Answer #1

Filling missing amount

Alpha Bravo Charlie
Sales 482000*6 = 2892000 420000 229000*2.5 = 572500
Net operating income 482000*24% = 115680 33600 40075
Average operating assets 482000 33600*100/24 = 140000 40075*100/17.5 = 229000
Margin 4% 8% 7%
Turnover 6 420000/140000 = 3 17.5/7 = 2.5
Return on investment 6*4 = 24% 24% 17.5%